# Recent Advances in Speech Language Models: A Survey

### v1

@@ 4.2.2 Language Model Instruction-Tuning
Several approaches have been proposed to construct instruction-following datasets for SpeechLMs. SpeechGPT Zhang et al. (2023a) and SpeechGPT-Gen Zhang et al. (2024) propose a two-stage instruction-tuning, including cross-modal instruction fine-tuning and chain-of-modality instruction fine-tuning. In the first stage, instruction data are generated based on ASR datasets by appending the instruction to paired ASR data, asking the model to convert speech into text. Similarly, paired data is also used to create instruction data for performing TTS. In the second stage, they construct a speech-in-speech-out dataset by transforming a text-based instruction-following dataset using TTS. Llama-Omni Fang et al. (2024) also creates instruction-following data by synthesizing text-based datasets, adhering to specific constraints. First, they transform the input text prompt into a format that mimics natural speech patterns. Next, they discard the original text response and employ a TextLM to generate answers to the converted prompts, ensuring these responses also follow natural speech patterns. Finally, they synthesize the prompt/response pairs using TTS. COSMIC Pan et al. (2023) constructed speech QA data by asking GPT-3.5 to generate question-answer pairs based on the transcriptions of English TED talk speeches. They showed the model trained on their proposed speech QA dataset can generalize to unseen tasks such as speech-to-text translation using in-context learning.
4.3 Speech Generation Paradigm
@@
In the previous sections, we discuss the typical generation paradigm for SpeechLMs, which involves taking a predefined input sequence and generating a complete response. However, this approach does not reflect the natural flow of voice interactions. For instance, during a conversation, one person may interrupt another, switching from listening to speaking. Additionally, a person might choose not to respond if the other is engaged in a conversation with someone else. Based on these observations, we identify two key aspects of advanced voice interaction skills for SpeechLMs: real-time interaction and silence mode.
Real-time Interaction refers to the capability of SpeechLMs to engage with users instantaneously. This interaction consists of two key components:
  1. 1.
    User Interruption: SpeechLMs should be able to be interrupted by users and should respond appropriately to new instructions provided during the conversation.
  2. 2.
    Simultaneous Response: SpeechLMs should be capable of generating responses while the user is still speaking.
@@
Both of these abilities require the model to effectively perform speech understanding (processing input) and speech generation (producing output) simultaneously. The study by dGSLM Nguyen et al. (2023b) introduces a dual-transformer architecture to model two-speaker dialogues, using one transformer to handle speech from each speaker. A cross-attention transformer layer is included to capture the interactions between the speakers’ content. In contrast, LSLM Ma et al. (2024) proposes a different approach, utilizing a single decoder-only Transformer to model one speaker’s speech in the dialogue. This model incorporates a streaming SSL encoder that continuously processes input from the listening channel and fuses its embeddings with those from the speaking channel.
@@
Silence Mode refers to the state in which the SpeechLMs remain inactive or silent during periods of non-interaction. This mode is essential for creating a natural conversational flow, allowing the model to avoid unnecessary interruptions. It is crucial for situations where a small group of users is having a discussion, as the SpeechLM needs to discern when to join in and when to stay silent. Additionally, it is important for the model to learn when to disregard instructions when users are not speaking at it. VITA Fu et al. (2024) is currently the only work that integrates silence mode. This method involves training the model on both query speech and non-query audio, which may include environmental sounds or non-query speech. As a result, the model learns to output the end-of-sequence token to terminate its response when non-query audio is detected.
5 Downstream Applications
@@
Table 4: Examples of the various capabilities of SpeechLMs.
Unlike traditional speech systems like ASR and TTS, which usually focus on specific tasks, SpeechLMs function as generative foundation models. They can handle a diverse array of speech-only, text-only, and multi-modal tasks by following various instructions. In this section, we explore the primary downstream applications of SpeechLMs. The tasks discussed here primarily consist of traditional speech-related tasks, along with some that are unique to SpeechLMs. In contrast to TextLMs, which generate text containing only semantic information, SpeechLMs can model both semantic and paralinguistic information—such as pitch and timbre—making them more powerful models. We thereby categorize the downstream applications of SpeechLMs into three main classes: semantic-related applications, speaker-related applications, and paralinguistic applications.
5.1 Semantic-Related Applications
@@
Keyword Spotting. Keyword spotting can be considered a special type of ASR, where its primary objective is to identify specific words or phrases within continuous speech. While traditional ASR systems aim to transcribe entire spoken utterances into text, keyword spotting focuses specifically on identifying and extracting predefined keywords or phrases within continuous speech. The primary application of keyword spotting is to build oice-activated assistants in smart home devices. Those devices are activated when the specific keywords are triggered. Therefore, although SpeechLMs are capable of spotting and understanding more than just a couple of words, keyword spotting can be used to efficiently trigger SpeechLMs to respond to user inputs.
@@
5.2 Speaker-Related Applications
@@
Speaker Diarization. Speaker diarization is the process of partitioning an audio stream into segments according to the identity of the speakers. It predicts “who is speaking when” for each timestamp Yang et al. (2021). A natural way to integrate speaker diarization into SpeechLMs is to have the model generate the transcript of each audio segment along with the identification of the speaker.
@@ 5.3 Paralinguistic Applications
6 Evaluations
Similar to TextLMs, SpeechLMs have a wide range of capabilities, making it challenging to compare different SpeechLMs. Consequently, it’s essential to evaluate SpeechLMs from various perspectives to determine their effectiveness. In this section, we review the commonly used methods and benchmarks for evaluating SpeechLMs. We categorize these evaluation methods into automatic and human assessments, each containing distinct evaluation aspects.
6.1 Automatic (Objective) Evaluation
@@
In the context of SpeechLMs, representation evaluation focuses on how well the model encodes speech features into meaningful vectors. GSLM Lakhotia et al. (2021) uses between-speaker ABX score to measure the embedding similarity. It quantifies how well-separated the phonetic categories are. Specifically, It works by comparing three sound samples: two from the same category (A) and one from a different category (B). The test measures how often the system correctly identifies that two sounds from category A are more similar to each other than one sound from A is to a sound from B. Another way of evaluating representations is through speech resynthesis Lakhotia et al. (2021). Specifically, an input speech is encoded into tokens and then synthesized back to speech. Then, word error rate (WER) or character error rate (CER) can be computed on the ASR results of the input and resynthesized speech. This measures the information loss caused by discretizing the input speech into speech tokens, thereby evaluating the robustness of the latent representations.
@@
Linguistic Evaluation. Linguistics, including lexical, syntactic, and semantic evaluation methods, assess the model’s ability to generate and understand the rules for constructing words, sentences, and meaningful contents. These evaluations focus on the correctness and appropriateness of word choices, the grammatical structure of the outputs, and the coherence and relevance of the generated content. In terms of benchmark datasets, sWUGGY Nguyen et al. (2020) assesses at the lexical level by determining if the model can distinguish a real word from a (real, non-real) word pair. sBLIMP Nguyen et al. (2020) evaluates at the syntactic level by determining if the model can identify the grammatically correct sentence from a (grammatical, ungrammatical) sentence pair. Spoken StoryCloze Hassid et al. (2024) evaluates semantic comprehension by assessing the model’s capability to select the genuine ending of a story from a pair of ending choices. All the evaluation is conducted by comparing the model’s negative log-likelihood of the data pair.
@@
Paralinguistic Evaluation. In contrast to linguistic evaluation, paralinguistic evaluation focuses on the non-verbal aspects of communication that accompany speech. Some works choose to utilize paralinguistic tokens alongside semantic tokens to enhance the paralinguistic abilities of SpeechLMs Kharitonov et al. (2021); Nguyen et al. (2024), so one way is to evaluate the paralinguistic tokens. pGSLM Kharitonov et al. (2021) measures the correctness, consistency, and expressiveness of the prosodic tokens. Correctness evaluates the model’s ability to generate accurate prosodic profiles by calculating the minimal mean absolute error (min-MAE) of the prosodic tokens from 20 generated samples against the prosodic tokens from the reference, consistency is assessed through the Pearson correlation between the mean values of the prompt prosodic and its generated continuation prosodic tokens, and expressiveness is measured by the standard deviation of the generated prosody token values, with the expectation that it matches the variability of the ground truth. We note that the same metrics can also be applied to other paralinguistic tokens. Instead of evaluating from the token level, Nguyen et al. (2024) propose to measure on the perceptual level. They introduced a speech-text sentiment preservation benchmark (STSP), which requires the model to generate a text or speech sequence of tokens that preserves the sentiment of the prompt. A sentiment classifier is used to assess the sentiment in the generated speech. It should be noted that although they only apply the preservation approach on sentiment, this idea can be generalized to other paralinguistic features, such as timbre or prosody.
@@
Generation Quality and Diversity. Quality and diversity are two crucial aspects of model generation. Typically, there is a trade-off between these dimensions when sampling model responses at different temperatures, so GSLM Lakhotia et al. (2021) suggests using the Area Under the Curve (AUC) with various temperature values. Specifically, AUC on perplexity and VERT are employed to assess these factors, where VERT represents the geometric mean of the ratio of k-grams in the generated speech that appears at least once. Additionally, the ChatGPT score can be utilized to evaluate the quality of the generated speech. In this process, the generated speech is transcribed using state-of-the-art ASR models and then sent to ChatGPT for quality (and diversity) assessment.
Downstream Evaluation. Downstream evaluation refers to evaluating the ability of SpeechLMs to perform specific tasks, such as ASR, TTS, Speaker Identification, etc. The evaluation can be performed on pre-trained models by adding few-shot example(s) at the start of the prompt or on the instruction-tuned models by directly instructing them to do so. SUPERB Yang et al. (2021) is a benchmark containing a wide range of downstream tasks that can be performed by SpeechLMs.
6.2 Human (Subjective) Evaluation.
@@
MOS is calculated by averaging the scores given by all evaluators for each audio sample, providing a single score that reflects the overall quality as perceived by humans. Variations of MOS focus on different aspects of speech quality, including MMOS, PMOS, and SMOS Kharitonov et al. (2021); Zhang et al. (2024). They evaluate the aspects of naturalness, prosody, and timbre similarity of the given speech, respectively.
@@
Typically, evaluating naturalness or timbre similarity involves collecting human opinions. However, this process can be complicated due to the challenges of recruiting participants and gathering their evaluations. As a result, researchers often turn to machine-based evaluations. They commonly employ neural network models specifically trained for these tasks. For instance, a naturalness prediction model Mittag et al. (2021) can assess the naturalness of generated outputs, while a speaker identification model can evaluate timbre similarity.
7 Challenges and Future Directions

### v2

@@ IV-B2 Language Model Instruction-Tuning
Several approaches have been proposed to construct instruction-following datasets for SpeechLMs. SpeechGPT [8] and SpeechGPT-Gen [46] propose a two-stage instruction-tuning, including cross-modal instruction fine-tuning and chain-of-modality instruction fine-tuning. In the first stage, instruction data are generated based on ASR datasets by appending the instruction to paired ASR data, asking the model to convert speech into text. Similarly, paired data is also used to create instruction data for performing TTS. In the second stage, they construct a speech-in-speech-out dataset by transforming a text-based instruction-following dataset using TTS. Llama-Omni [11] also creates instruction-following data by synthesizing text-based datasets, adhering to specific constraints. First, they transform the input text prompt into a format that mimics natural speech patterns. Next, they discard the original text response and employ a TextLM to generate answers to the converted prompts, ensuring these responses also follow natural speech patterns. Finally, they synthesize the prompt/response pairs using TTS. COSMIC [54] constructed speech QA data by asking GPT-3.5 to generate question-answer pairs based on the transcriptions of English TED talk speeches. They showed the model trained on their proposed speech QA dataset can generalize to unseen tasks such as speech-to-text translation using in-context learning.
IV-B3 Language Model Post-Alignment
Post-alignment is the critical process of refining a language model’s behavior to align with human preferences, ensuring that its outputs are both safe and reliable. This stage is typically regarded as the final phase of language model training. It often employs techniques like Reinforcement Learning from Human Feedback (RLHF), specifically methods such as Proximal Policy Optimization (PPO) [118] and Direct Preference Optimization (DPO) [119].
Post-alignment in SpeechLMs focuses on addressing the unique challenges inherent in the speech interaction pipeline. Align-SLM [120] identifies that SpeechLMs often generate inconsistent semantic content when given the same prompt. It tackles this by using a TextLM to choose the preferred response from SpeechLMs after transcribing them via ASR and then aligning those preferences using DPO. On the other hand, SpeechAlign [121] concentrates on the acoustic quality of SpeechLMs. It observes that differences between the “golden” speech tokens and those generated by the language model lead to subpar acoustic quality in the generated speech as the vocoder synthesizes speech from generated tokens during inference. To mitigate this, it employs various optimization techniques to align the language model’s output with the distribution of the “golden” tokens. Despite its importance, the post-alignment of SpeechLMs remains under-explored. A critical application of post-alignment is to mitigate the safety risks associated with generative models. Thus, future research should prioritize identifying and addressing the unique safety challenges posed by SpeechLMs (see Section VII-D).
IV-C Speech Generation Paradigm
@@
Most approaches covered in earlier sections follow the traditional generation paradigm of SpeechLMs, which involves taking a predefined input sequence and generating a complete response. However, this approach does not reflect the natural flow of voice interactions. For instance, during a conversation, one person may interrupt another, switching from listening to speaking. Additionally, a person might choose not to respond if the other is engaged in a conversation with someone else. Based on these observations, we identify two key aspects of advanced voice interaction skills for SpeechLMs: real-time interaction and interactive period recognition.
@@
Real-time Interaction refers to the capability of SpeechLMs to engage with users instantaneously. This interaction consists of two key components. 1) User Interruption: SpeechLMs should be able to be interrupted by users and should respond appropriately to new instructions provided during the conversation. 2) Simultaneous Response: SpeechLMs should be capable of generating responses while the user is still speaking. This requires the model to effectively perform speech understanding (processing input) and speech generation (producing output) simultaneously. Some literature refers to the ability as full-duplex modeling. dGSLM [4] introduces a dual-transformer architecture to model two-speaker dialogues, using one transformer to handle speech from each speaker. A cross-attention transformer layer is included to capture the interactions between the speakers’ content. However, most approaches rely on a single language model architecture to manage both channels in a dialogue. Parrot [45] introduces the ”next-token-pair prediction” approach, which uses a single decoder-only transformer to model both channels simultaneously by predicting tokens for both channels at each time step. Moshi [9] concatenates the user input channel and the model response channel and employs an RQ-Transformer that first processes data along the temporal dimension and then along the channel dimension. LSLM [51] utilizes a decoder-only Transformer focusing on modeling one speaker’s speech in the dialogue. This model incorporates a streaming SSL encoder that continuously processes input from the listening channel and fuses its embeddings with those from the speaking channel.
Interactive Period Recognition (IPR) refers to the ability of SpeechLMs to discern whether users are actively interacting with them or not. During interactive periods, the model provides responses, while during non-interactive periods, it remains silent. IPR is essential for maintaining a natural conversational flow, preventing unnecessary interruptions, and ensuring the model only participates when appropriate. It is crucial for situations where a small group of users is having a discussion, as the SpeechLM needs to discern when to join in and when to stay silent. Additionally, it is important for the model to learn when to disregard instructions when users are not speaking at it.
@@
Interactive Period Recognition (IPR) refers to the ability to recognize whether the users are interacting with it or not. SpeechLMs should provide response during the interactive period and remain silent during the non-interactive period. IPR is essential for creating a natural conversational flow, allowing the model to avoid unnecessary interruptions. It is crucial for situations where a small group of users is having a discussion, as the SpeechLM needs to discern when to join in and when to stay silent. Additionally, it is important for the model to learn when to disregard instructions when users are not speaking at it. One approach to achieving IPR is through a Voice Activity Detection (VAD) module. MiniCPM-o 2.6 [122] integrates a VAD module to ensure the model responds only when the input audio surpasses a predefined VAD threshold. Inputs below this threshold are considered noise and ignored. VITA [55] takes a different approach by training the SpeechLM to distinguish between query speech and non-query audio. The model learns to output an end-of-sequence token to terminate its response when non-query audio is detected.
V Downstream Applications
TABLE V: Examples of the various capabilities of SpeechLMs.
@@ V Downstream Applications
Unlike traditional speech systems like ASR and TTS, which usually focus on specific tasks, SpeechLMs function as generative foundation models. They can handle a diverse array of speech-only, text-only, and multi-modal tasks by following various instructions. In this section, we explore the primary downstream applications of SpeechLMs. The tasks discussed here primarily consist of traditional speech-related tasks, along with some that are unique to SpeechLMs. In contrast to TextLMs, which generate text containing only semantic information, SpeechLMs can model both semantic and paralinguistic information—such as pitch and timbre—making them more powerful models. We thereby categorize the downstream applications of SpeechLMs into three main classes: semantic-related applications, speaker-related applications, and paralinguistic applications. Table V provides an example for each downstream task.
V-A Semantic-Related Applications
@@
Keyword Spotting. Keyword spotting can be considered a special type of ASR, where its primary objective is to identify specific words or phrases within continuous speech. While traditional ASR systems aim to transcribe entire spoken utterances into text, keyword spotting focuses specifically on identifying and extracting predefined keywords or phrases within continuous speech. The primary application of keyword spotting is to build voice-activated assistants in smart home devices. Those devices are activated when the specific keywords are triggered. Therefore, although SpeechLMs are capable of spotting and understanding more than just a couple of words, keyword spotting can be used to efficiently trigger SpeechLMs to respond to user inputs.
@@
V-B Speaker-Related Applications
@@
Speaker Diarization. Speaker diarization is the process of partitioning an audio stream into segments according to the identity of the speakers. It predicts “who is speaking when” for each timestamp [123]. A natural way to integrate speaker diarization into SpeechLMs is to have the model generate the transcript of each audio segment along with the identification of the speaker.
@@ V-C Paralinguistic Applications
VI Evaluations
TABLE VI: A summary of popular benchmarks for the evaluation of SpeechLMs. I/O, $A$, and $T$ represent input/output format, audio, and text, respectively.
Name Eval Type # Tasks Audio Type I/O
ABX [124, 125, 126] Representation 1 Speech $A\rightarrow-$
sWUGGY [126] Linguistic 1 Speech $A\rightarrow-$
sBLIMP [126] Linguistic 1 Speech $A\rightarrow-$
sStoryCloze [38] Linguistic 1 Speech $A/T\rightarrow-$
STSP [5] Paralinguistic 1 Speech $A/T\rightarrow A/T$
MMAU [127] Downstream 27 Speech, Sound, Music $A\rightarrow T$
Audiobench [128] Downstream 8 Speech, Sound $A\rightarrow T$
AIR-Bench [129] Downstream 20 Speech, Sound, Music $A\rightarrow T$
SD-Eval [130] Downstream 4 Speech $A\rightarrow T$
SUPERB [131] Downstream 10 Speech $A\rightarrow T$
Dynamic-SUPERB [131] Downstream 180 Speech, Sound, Music $A\rightarrow T$
SALMON [132] Downstream 8 Speech $A\rightarrow-$
VoiceBench [133] Downstream 8 Speech $A\rightarrow A$
VoxEval [134] Downstream 56 Speech $A\rightarrow A$
Similar to TextLMs, SpeechLMs have a wide range of capabilities, making it challenging to compare different SpeechLMs. Consequently, it’s essential to evaluate SpeechLMs from various perspectives to determine their effectiveness. In this section, we review the commonly used methods and benchmarks (Table VI) for evaluating SpeechLMs. We categorize these evaluation methods into automatic and human assessments, each containing distinct evaluation aspects.
VI-A Automatic (Objective) Evaluation
@@
In the context of SpeechLMs, representation evaluation focuses on how well the model encodes speech features into meaningful vectors. GSLM [37] uses between-speaker ABX score to measure the embedding similarity. It quantifies how well-separated the phonetic categories are. Specifically, It works by comparing three sound samples: two from the same category (A) and one from a different category (B). The test measures how often the system correctly identifies that two sounds from category A are more similar to each other than one sound from A is to a sound from B. Another way of evaluating representations is through speech resynthesis [37]. Specifically, an input speech is encoded into tokens and then synthesized back to speech. Then, word error rate (WER) or character error rate (CER) can be computed on the ASR results of the input and resynthesized speech. This measures the information loss caused by discretizing the input speech into speech tokens, thereby evaluating the robustness of the latent representations.
@@
Linguistic Evaluation. Linguistics, including lexical, syntactic, and semantic evaluation methods, assess the model’s ability to generate and understand the rules for constructing words, sentences, and meaningful contents. These evaluations focus on the correctness and appropriateness of word choices, the grammatical structure of the outputs, and the coherence and relevance of the generated content. In terms of benchmark datasets, sWUGGY [126] assesses at the lexical level by determining if the model can distinguish a real word from a (real, non-real) word pair. sBLIMP [126] evaluates at the syntactic level by determining if the model can identify the grammatically correct sentence from a (grammatical, ungrammatical) sentence pair. Spoken StoryCloze [38] evaluates semantic comprehension by assessing the model’s capability to select the genuine ending of a story from a pair of ending choices. All the evaluation is conducted by comparing the model’s negative log-likelihood of the data pair.
@@
Paralinguistic Evaluation. In contrast to linguistic evaluation, paralinguistic evaluation focuses on the non-verbal aspects of communication that accompany speech. Some works choose to utilize paralinguistic tokens alongside semantic tokens to enhance the paralinguistic abilities of SpeechLMs [42, 5], so one way is to evaluate the paralinguistic tokens. pGSLM [42] measures the correctness, consistency, and expressiveness of the prosodic tokens. Correctness evaluates the model’s ability to generate accurate prosodic profiles by calculating the minimal mean absolute error (min-MAE) of the prosodic tokens from 20 generated samples against the prosodic tokens from the reference, consistency is assessed through the Pearson correlation between the mean values of the prompt prosodic and its generated continuation prosodic tokens, and expressiveness is measured by the standard deviation of the generated prosody token values, with the expectation that it matches the variability of the ground truth. We note that the same metrics can also be applied to other paralinguistic tokens. Instead of evaluating from the token level, SPIRIT-LM [5] proposes to measure on the perceptual level. They introduced a speech-text sentiment preservation benchmark (STSP), which requires the model to generate a text or speech sequence of tokens that preserves the sentiment of the prompt. A sentiment classifier is used to assess the sentiment in the generated speech. It should be noted that although they only apply the preservation approach on sentiment, this idea can be generalized to other paralinguistic features, such as timbre or prosody.
@@
Generation Quality and Diversity. Quality and diversity are two crucial aspects of model generation. Typically, there is a trade-off between these dimensions when sampling model responses at different temperatures, so GSLM [37] suggests using the Area Under the Curve (AUC) with various temperature values. Specifically, AUC on perplexity and VERT are employed to assess these factors, where VERT represents the geometric mean of the ratio of k-grams in the generated speech that appears at least once. Additionally, the ChatGPT score can be utilized to evaluate the quality of the generated speech. In this process, the generated speech is transcribed using state-of-the-art ASR models and then sent to ChatGPT for quality (and diversity) assessment.
Real-Time Interaction Evaluation. Real-time interaction evaluation involves assessing the ability of SpeechLMs to interact in real-time, which is crucial for those models that facilitate full-duplex communication. Current research focuses on evaluating the naturalness and usefulness of speech generated in real-time. dGSLM [4] examines the naturalness of dialogues between two speakers by introducing different turn-taking events, such as speech segments (Inter-Pausal Unit, IPU), pauses within speech (pause), pauses between speeches (gaps), and overlapping speech (overlap). The generated speech is considered more natural if the statistics of these turn-taking events closely resemble those found in human dialogues. Another method involves speech continuation, where the generated speech is considered more natural if the turn-taking statistics of the speech prompts align closely with those of the subsequent continuations. Additionally, evaluating SpeechLM’s usefulness as an AI assistant in real-time interaction is crucial. The Parrot model [45] suggests using reflective pauses and interruptions to assess usefulness, where reflective pause evaluates the SpeechLM’s ability to remain silent while the user speaks, and interruption measures the SpeechLM’s capability to cease speaking when interrupted.
Downstream Evaluation. Downstream evaluation refers to evaluating the ability of SpeechLMs to perform specific tasks, such as ASR, TTS, Speaker Identification, etc. The evaluation can be performed on pre-trained models by adding few-shot example(s) at the start of the prompt or on the instruction-tuned models by directly instructing them to do so. Several benchmarks compile a broad array of downstream tasks to provide a comprehensive assessment of SpeechLMs. SUPERB [123] includes a variety of speech understanding tasks. SD-Eval [130] assesses the paralinguistic understanding of SpeechLMs using the emotion, age, environment, and age classification tasks. SALMON [132] tests the SpeechLMs’ ability to generate speech with consistent paralinguistic and environmental characteristics. Dynamic-SUPERB [131], MMAU [127], AirBench [129], and AudioBench [128] extend beyond traditional speech tasks to include sound and/or music-related challenges as well.
While these benchmarks offer comprehensive coverage of audio-related tasks, they primarily require models to respond in text, which creates a barrier to end-to-end speech interaction evaluation. To address this limitation, newer benchmarks have emerged that support speech output. VoxEval [134] focuses on benchmarking the knowledge understanding capabilities of SpeechLMs, providing question-answer pairs in comprehensive subjects in speech format, along with an evaluation pipeline tailored for speech output. Additionally, it presents questions under various input audio conditions for evaluating the model’s robustness. Similarly, VoiceBench [133] supports speech output while focusing on assessing general knowledge, instruction following, and response safety of SpeechLMs.
VI-B Human (Subjective) Evaluation.
@@
MOS is calculated by averaging the scores given by all evaluators for each audio sample, providing a single score that reflects the overall quality as perceived by humans. Variations of MOS focus on different aspects of speech quality, including MMOS, PMOS, and SMOS [42, 46]. They evaluate the aspects of naturalness, prosody, and timbre similarity of the given speech, respectively.
@@
Typically, evaluating naturalness or timbre similarity involves collecting human opinions. However, this process can be complicated due to the challenges of recruiting participants and gathering their evaluations. As a result, researchers often turn to machine-based evaluations. They commonly employ neural network models specifically trained for these tasks. For instance, a naturalness prediction model [135] can assess the naturalness of generated outputs, while a speaker identification model can evaluate timbre similarity.
VII Challenges and Future Directions